OpenStudy (anonymous):

How can I find sixth term of a geometric sequence for which A(subscript)2=288 and r=1/4?

OpenStudy (cwrw238):

the formula for the nth term is a*r^(n-1) where a = first term and r = common ratio

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Okay

OpenStudy (cwrw238):

you can find the first term by the relation second term / first term = common ratio

OpenStudy (cwrw238):

second term ( A2) = 288

OpenStudy (anonymous):

how would I find the first term?

OpenStudy (cwrw238):

secomd term = 288 common ratio = 1/4 A2 / A1 = r 288/A1 = 1/4 find A1